The measure of angle A is 6° greater than the measure of angleB. The two angles are complementary. Find the measure of each angle. I don't get this question can you please help me.
Answer by Alan3354(69443)

The measure of angle A is 6° greater than the measure of angleB. The two angles are complementary. Find the measure of each angle. I don't get this question can you please help me.
---------
The measure of angle A is 6° greater than the measure of angle B
--> A = B + 6
==========
The two angles are complementary.
--> A + B = 90
---------------------
A = B + 6
-------------------
A - B = 6
A + B = 90
------------------------------- Add
zat make sense?